Two objects attract each other with a gravitational force of magnitude 9.40 10-9 N when separated by 19.8 cm. If the total mass of the objects is 5.03 kg, what is the mass of each?

heavier mass:

lighter mass:
Answer #1

Let the mass of Heavier be mh​​​​​, and the mass of lighter be ml

Given, mh + ml = 5.03 kg [Eq1]

F = 9.40*10^-9 N

r = 19.8 cm

By using, F = Gmhml/r^2

9.40*10^-9 = (6.674*10^-11)mhml/(0.198)^2

mh ml = 9.4*10^-9*(0.198)^2 / (6.674*10^-11) = 5.52

From Eq1,

(5.03 - ml​​​​​​) *ml = 5.52

5.03ml - ml^2 = 5.52

ml^2 - 5.03ml + 5.52 = 0

ml = 1.62

mh = 5.03 - 1.62 = 3.41

m​​​​